These were accepted as a reasonable account of the last meeting.
Alastair will check that script was included in the stable release.
All pages with review date prior to August 2015 should be reviewed by the end of April.
Neil has stopped this for now. The logs are being sent to more than one place. Neil will fix.
Tapes to be wiped.
Gordon will send out email to sys-announce.
The DNS has been delegated to us so there is now a dns/bayes file and appropriate web forwarding in place.
Blog articles suggested:
Report from Computing Executive Group
We have been given permission from College to appoint an additional CSO.
Reports from Units
There are 2 KVM servers in AT now - oyster and waterloo. They are available for use. All 3 sites now have multiple supported KVM servers between which VMs can be migrated.
US's wiki page is here.
Topics for discussion
The next meeting will be on Wednesday 26th April at 10:00, in IF-4.31, chaired by Craig.
Please contact us with any comments or corrections.
Unless explicitly stated otherwise, all material is copyright The University of Edinburgh